Susan wants to invest 5000. If she invests part at 7% simple interest, part at 6%, and receives a total of 332 after one year, how much does she invest at each rate? I need help with the formula and solution, please.
Answer by checkley77(12569)

.07X+.06(5,000-X)=332
.07X+300-.06X=332
.01X=332-300
.01X=32
X=32/.01
X=3,200 AMOUNT INVESTED $ 7%.
5,000-3,200=1,800 INVESTED @ 6%.
PROOF:
.07*3,200+.06*1,800=332
224+108=332
332=332